### v4

@@ 3 Preliminaries
In this section, we first introduce the background of feature embeddings, which is fundamental for most feature interaction models based on deep learning. To help understand our proposed model GraphFM, which are based on FMs and GNNs, we then describe these two lines of work.
@@ 3.1 Feature Embeddings
In many real-world predictive tasks, such as CTR prediction, input instances consist of both sparse categorical and numerical features. By tradition, we represent each input instance as a sparse vector:
@@
where $n$ is the number of feature fields and $\mathbf{x}_{i}$ is the representation of the $i$-th feature field (aka feature). Since the categorical features are very sparse and high-dimensional, a common way is to map them into a low-dimensional latent space. Specifically, a categorical feature $\mathbf{x}_{i}$ is mapped to dense embedding $\mathbf{e}_{i}\in\mathbb{R}^{d}$ as:
@@ 3.2 Factorization Machines
Factorization machine (FM) is originally proposed for collaborative recommendation Rendle (2010, 2012). It estimates the target by modeling all interactions between each pair of features:
@@
In principle, FMs can be extended to higher-order feature combinations Rendle (2010, 2012). Let $k\in\{2,\dots,K\}$ denote the order or degree of feature interactions considered and $\textbf{e}_{i}^{(k)}$ denotes the embedding of feature $i$ for order $k$, the $K$-order higher-order FM (HOFM) can be defined as
@@
where $\langle\textbf{e}_{i_{1}}^{(K)},\dots,\textbf{e}_{i_{K}}^{(K)}\rangle=\text{sum}(\textbf{e}_{i_{1}}^{(K)}\odot\dots\odot\textbf{e}_{i_{K}}^{(K)})$ (sum of element-wise products). Since all the feature interactions of order up to $K$ are included, its time complexity increases exponentially with, resulting in high computational complexity. Considering that not all feature interactions are beneficial, FMs have trouble modeling higher-order feature interactions in terms of both efficiency and effectiveness.
Although several recent researches have enhanced FMs with DNNs to model higher-order feature interactions, like NFM He and Chua (2017) and DeepFM Guo et al. (2017), they model the higher-order feature interaction in an implicit manner, which lack persuasive rationales on the model outcome.
@@ 3.3 Graph Neural Networks
Given a graph $G=\left\{V,E\right\}$ denote a graph, GNNs learn the representation vectors of nodes by exploring their correlations with neighboring nodes. Modern GNNs follow a neighborhood aggregation strategy, where we iteratively update the representation of a node by aggregating the feature of its neighbors Xu et al. (2018). After $k$ iterations of aggregation, a node’s representation encodes its interaction with neighbors within $k$ hops. The choice of aggregation strategy in GNNs is crucial. A number of architectures have been proposed. In this work, we adopt the attentional aggregation strategy in Veličković et al. (2018), which will be further elaborated in Section 4.
@@ 4.1 Model Overview
The overview of GraphFM is shown in Fig. 1. Each input instance (multi-field feature) is represented as a graph where the nodes are feature fields, and the edges are interactions Li et al. (2019); Su et al. (2020). Note that we will use nodes and features, edges, and interactions interchangeably in this paper. GraphFM updates feature representations layer by layer. The feature embeddings described in Section 3.1 are taken as the initial feature embeddings of GraphFM, i.e., $\textbf{e}^{(1)}_{i}=\textbf{e}_{i}$, where $\mathbf{e}^{(k)}_{i}$ stands for the updated feature embeddings at $k$-th layer. Since no edge information is given, we need to select the edges (beneficial interactions) by the interaction selection component first. Then we aggregate these selected feature interactions to update feature embeddings in the neighborhood aggregation component. Within each $k$-th layer, we are able to select and model only the beneficial $k$-th order feature interactions and encode these factorized interactions into feature representations. Finally, these learned feature embeddings encoded interactions of order up to $K$ are concatenated to make the final prediction.
There are mainly two components in each layer of GraphFM. Next, we will introduce them in detail. As we focus on describing the detailed mechanism at every single layer, we omit the layer index $k$ if not necessary.
Figure 1: The overview of GraphFM. The input features are modeled as a graph, where nodes are feature fields, and edges are interactions. At each layer of GraphFM, the edges (beneficial interactions) are first selected by the interaction selection component. Then these selected feature interactions are aggregated via the attention net to update feature embeddings in the interaction aggregation component. The learned feature embeddings at every layer are used for the final prediction jointly.
@@ 4.2 Interaction Selection
To select the beneficial pairwise feature interactions, we devised the interaction selection mechanism. This can also be viewed as inferring the graph structure, which is to predict the links between nodes. However, the graph structure $G=\left\{V,E\right\}$ is discrete, where an edge $(v_{i},v_{j})\in E$ linking two nodes is either present or absent. This makes the process non-differentiable, therefore it cannot be directly optimized with gradient-descent-based optimization techniques.
To overcome this limitation, we replace the edge set $E$ with weighted adjacency $\mathbf{P}$, where $p_{ij}$ is interpreted as the probability of $(v_{i},v_{j})\in E$, which also reflects how beneficial their interaction is. It should be noted that we learn different graph structures $\mathbf{P}^{(k)}$ at each $k$-th layer. Compared with using a fixed graph at each layer, we have more efficiency and flexibility in enumerating the beneficial higher-order feature interaction by this means. More specifically, by using a fixed graph structure at each layer, we can only obtain a fixed set of feature interactions. However, with the adaptive learned graph structure at each layer, the model is capable of modeling any potential feature interactions.
@@ 4.2.1 Metric function
We aim to design a metric function between each pair of feature interactions to measure whether they are beneficial. Formally, the weight $p_{ij}$ of an edge $(v_{i},v_{j})$ is computed via a metric function $f_{s}(\mathbf{e}_{i},\mathbf{e}_{j})$. We here adopt a Neural Matrix Factorization (NMF) He et al. (2017) based function to estimate the edge weight. Formally, a Multi-Layer Perception (MLP) with one hidden layer is used to transform the element-wise product of these feature vectors to a scalar:
@@
where $\mathbf{W}^{s}_{1}$, $\mathbf{W}^{s}_{2}$, $\mathbf{b}^{s}_{1}$, $\mathbf{b}^{s}_{2}$ are the parameters of MLP. $\delta(\cdot)$ and $\sigma(\cdot)$ are ReLU and sigmoid activation functions, respectively. It should be noted that $f_{s}$ is invariant to the order of its input, i.e., $f_{s}(\mathbf{e}_{i},\mathbf{e}_{j})=f_{s}(\mathbf{e}_{j},\mathbf{e}_{i})$. Therefore, the estimated edge weights are identical to the same pair of nodes. Such continuous modeling of graph structure enables backpropagation of the gradients. Since we do not have a ground-truth graph structure, the gradients come from the errors between the model output with the target.
Intuitively, we treat the element-wise product of each pair of feature embeddings as a term and estimate its weight using MLP. One can also choose the Euclidean distance Kazi et al. (2020) or other distance metrics.
@@ 4.2.2 Graph sampling
From the estimated edge weighted matrix $\mathbf{P}^{(k)}$ at each layer, we then sample the beneficial feature interactions, which is also to sample the neighborhood for each feature field. In this work, we uniformly sample a fixed-size set of neighbors. For each feature node $v_{i}$ at $k$-th layer, we select $m_{k}$ edges according to the first $m_{k}$ elements of $\mathbf{P}^{(k)}[i,:]$, which can be illustrated as follows:
@@
where $\mathbf{P}^{(k)}[i,:]$ denotes the $i$-th column of matrix $\mathbf{P}^{(k)}$ at $k$-th layer, $\mathbf{P}^{(k)}[i,-\text{idx}_{i}]$ contains a subset of columns of $\mathbf{P}^{(k)}$ that are not indexed by $\text{idx}_{i}$. $\text{argtop}_{m_{k}}$ is an operator that performs the $m_{k}$-most important nodes selection for the query node $i$ to attend. We only keep these $m_{k}$ feature nodes, and the others are masked. Thus the neighborhood set of node $v_{j}$ is defined as:
@@
Practically speaking, we found that our approach could achieve high performance when $k=3$, $m_{1}$ equals the number of feature fields, which means that in the first layer, we model all pairs of feature interactions.
It is worth mentioning that we have also tried to set a threshold to select the edges in the graph, i.e., setting a minimum value for the edge probability of cutting edges off. But the performance is not as good as using a fixed-degree graph. This is reasonable as the edge weights of different nodes’ neighbors are at different scales. Setting a single threshold on all the nodes will lead to the situation that the numbers of nodes’ neighbors vary a lot. Some nodes will have barely any adjacent nodes after cutting off, while some may still have many.
@@ 4.3 Interaction Aggregation
Since we have selected the beneficial feature interactions, or in other words, learned the graph structure, we perform the interaction (neighborhood) aggregation operation to update feature representations.
@@
This indicates the importance of the interactions between feature $v_{i}$ and features $v_{j}$.
Note that we only compute $c_{ij}$ for nodes $j\in\mathcal{N}_{i}$, where $\mathcal{N}_{i}$ denotes the neighborhood of node $v_{i}$, which is also the set of features whose interaction with $v_{i}$ are beneficial. To make coefficients easily comparable across different feature nodes, we normalized them across all choices of $j$ using a softmax function:
@@
Once obtained the normalized attention coefficients, we compute the linear combination of these feature interactions with a nonlinearity as the updated feature representations:
@@
where $\alpha_{ij}$ measures the attention coefficients of each feature interaction between feature $i$ and $j$, while $p_{ij}$ stands for the probability of this feature interaction being beneficial. The attention coefficient $\alpha_{ij}$ is calculated by the soft attention mechanism, while the $p_{ij}$ is calculated by the hard attention mechanism. By multiplying them together, we control the information of selected feature interactions and make the parameters in the interaction selection component trainable with gradient back-propagation.
To capture the diversified polysemy of feature interactions in different semantic subspaces Li et al. (2020) and also stabilize the learning process Vaswani et al. (2017); Veličković et al. (2018), we extend our mechanism to employ multi-head attention. Specifically, $H$ independent attention mechanisms execute the update of Equation 12, and then these features are concatenated, resulting in the following output feature representation:
@@
where $\parallel$ denotes concatenation, $\alpha_{ij}^{h}$ is the normalized attention coefficient computed by the $h$-th attention mechanism, and $\mathbf{W}^{h}_{a}$ is the corresponding linear transformation matrix. One can also choose to employ average pooling to update the feature representations:
@@ 4.4 Prediction and Optimization
The output of each $k$-th layer, is a set of $n$ feature representation vectors encoded feature interactions of order up to $k$, namely $\left\{\mathbf{e}^{(k)}_{1},\mathbf{e}^{(k)}_{2},\dots,\mathbf{e}^{(k)}_{n}\right\}$. Since the representations obtained in different layers encode the interactions of different orders, they have different contributions to the final prediction. As such, we concatenate them to constitute the final representation of each feature Wang et al. (2019):
@@
At last, We employ average pooling on the vectors of all features to obtain a graph-level output and use a projection vector $\mathbf{p}$ to make the final prediction:
@@
where $N$ is the total number of training instances, and $\sigma$ denotes the sigmoid function. $y_{i}$ and $\hat{y}_{i}$ denote the label of instance $i$ and the prediction of GraphFM for it, respectively. The model parameters are updated using AdamKingma and Ba (2014).
@@ 5 Experiments
Table 1: Statistics of evaluation datasets.
@@ 5.1 Experimental Settings
Our experiments are conducted on three real-world datasets, two CTR benchmark datasets, and one recommender system dataset. Details of these datasets are illustrated in Table 1. The data preparation follows the strategy in Tian et al. (2023). We randomly split all the instances in 8:1:1 for training, validation, and testing. We adopt the two most popular metrics, AUC and Logloss to measure the probability that one prediction diverges from the ground truth.
@@ 5.1.1 Baselines
We compare GraphFM with four classes of state-of-the-art methods:(A) the linear approach that only uses individual features; (B) FM-based methods that consider second-order feature interactions; (C) DNN-based methods that model high-order feature interactions; (D) aggregation-based methods that update features’ representation and model their high-order feature interactions via aggregation strategy.
@@
  • LR (A) refers to the linear/logistics regression, which can only model linear interactions.
  • FM Rendle (2012)  (B) is the official implementation for FM, which can only model second-order interactions.
  • AFM Xiao et al. (2017) (B) is an extension of FM, which considers the weights of different second-order feature interactions by using attention mechanisms.
  • AFN Cheng et al. (2020)  (B) learns arbitrary-order feature interactions adaptively from data, instead of explicitly modeling all the cross features within a fixed maximum order.
@@
  • NFM He and Chua (2017) (C) devises a bi-Interaction layer to model second-order interactions and uses DNN to introduce non-linearity and model high-order interactions.
@@
  • DeepCrosssing Shan et al. (2016)  (C) utilizes DNN with residual connections to model non-linear feature interactions in an implicit manner.
  • CrossNet Wang et al. (2017) (C) is the core of Deep&Cross model, which models feature interactions explicitly by taking the outer product of concatenated feature vector at the bit-wise level.
  • xDeepFM Lian et al. (2018)  (C) takes the outer product of the stacked feature matrix at a vector-wise level to model feature interactions explicitly. It can be also combined with DNN which models implicit and explicit interactions simultaneously.
  • DCNV2 Wang et al. (2021)  (C) utilizes cross network from DCN to learning explicit and bounded-degree cross features.
  • AutoInt Song et al. (2019) (D) uses self-attentive network to learn high-order feature interactions explicitly. It can also be seen as performing the multi-head attention mechanism Vaswani et al. (2017) on the fully-connected graph.
@@
  • InterHAt Li et al. (2020) (D) utilizes an attention mechanism to aggregate features, which will then be multiplied by the original features to produce the higher-order feature interactions.

### v5

@@ 3 Preliminaries
In this section, we first introduce the background of feature embeddings, which are fundamental for most feature interaction models based on deep learning. To help understand our proposed model GraphFM, which is based on FMs and GNNs, we then describe these two lines of work.
@@ 3.1 Feature Embeddings
In many real-world predictive tasks, such as CTR prediction, input instances consist of both sparse categorical and numerical features. Traditionally, we represent each input instance as a sparse vector:
@@
where $n$ is the number of feature fields and $\mathbf{x}_{i}$ is the representation of the $i$-th feature field (aka feature). Since categorical features are very sparse and high-dimensional, a common way is to map them into a low-dimensional latent space. Specifically, a categorical feature $\mathbf{x}_{i}$ is mapped to dense embedding $\mathbf{e}_{i}\in\mathbb{R}^{d}$ as:
@@ 3.2 Factorization Machines
The factorization machine (FM) was originally proposed for collaborative recommendation Rendle (2010, 2012). It estimates the target by modelling all interactions between each pair of features:
@@
In principle, FMs can be extended to higher-order feature combinations Rendle (2010, 2012). Let $k\in\{2,\dots,K\}$ denote the order or degree of feature interactions considered and $\textbf{e}_{i}^{(k)}$ denote the embedding of feature $i$ for order $k$, the $K$-order higher-order FM (HOFM) can be defined as
@@
where $\langle\textbf{e}_{i_{1}}^{(K)},\dots,\textbf{e}_{i_{K}}^{(K)}\rangle=\text{sum}(\textbf{e}_{i_{1}}^{(K)}\odot\dots\odot\textbf{e}_{i_{K}}^{(K)})$ (sum of element-wise products). Since all the feature interactions of order up to $K$ are included, its time complexity increases exponentially, resulting in high computational complexity. Considering that not all feature interactions are beneficial, FMs have trouble modelling higher-order feature interactions in terms of both efficiency and effectiveness.
Although several recent studies have enhanced FMs with DNNs to model higher-order feature interactions, like NFM He and Chua (2017) and DeepFM Guo et al. (2017), they model higher-order feature interactions in an implicit manner, and lack persuasive rationales for model outcomes.
@@ 3.3 Graph Neural Networks
Given that a graph $G=\left\{V,E\right\}$ denotes a graph, GNNs learn the representation vectors of nodes by exploring their correlations with neighboring nodes. The modified GNNs follow a neighborhood aggregation strategy, where we iteratively update the representation of a node by aggregating the features of its neighbors Xu et al. (2018). After $k$ iterations of aggregation, a node’s representation encodes its interaction with neighbors within $k$ hops. The choice of aggregation strategy for GNNs is crucial. A number of architectures have been proposed. In this work, we adopt the attentional aggregation strategy in Veličković et al. (2018), which will be further elaborated upon in Section 4.
@@ 4.1 Model Overview
An overview of GraphFM is shown in Fig. 1. Each input instance (multi-field feature) is represented as a graph where the nodes are feature fields, and the edges are interactions Li et al. (2019); Su et al. (2020). Note that we use nodes and features, edges, and interactions interchangeably in this paper. GraphFM updates feature representations layer by layer. The feature embeddings described in Section 3.1 are taken as the initial feature embeddings of GraphFM, i.e., $\textbf{e}^{(1)}_{i}=\textbf{e}_{i}$, where $\mathbf{e}^{(k)}_{i}$ represents the updated feature embeddings at the $k$-th layer. Since no edge information is given, we need to select the edges (beneficial interactions) by the interaction selection component first. Then, we aggregate these selected feature interactions to update the feature embeddings in the neighborhood aggregation component. Within each $k$-th layer, we are able to select and model only the beneficial $k$-th order feature interactions and encode these factorized interactions into feature representations. Finally, these learned feature embedding encoded interactions of order up to $K$ are concatenated to make the final prediction.
There are two main components in each layer of GraphFM. Next, we will introduce them in detail. As we focus on describing the detailed mechanism at every single layer, we omit the layer index $k$ if not necessary.
Figure 1: Overview of GraphFM. The input features are modelled as a graph, where nodes are feature fields, and edges are interactions. At each layer of GraphFM, the edges (beneficial interactions) are first selected by the interaction selection component. Then these selected feature interactions are aggregated via the attention network to update the feature embeddings in the interaction aggregation component. The learned feature embeddings at every layer are used for the final prediction jointly.
@@ 4.2 Interaction Selection
To select beneficial pairwise feature interactions, we devised an interaction selection mechanism. This can also be viewed as inferring the graph structure, which predicts the links between nodes. However, the graph structure $G=\left\{V,E\right\}$ is discrete, where an edge $(v_{i},v_{j})\in E$ linking two nodes is either present or absent. This makes the process non-differentiable; therefore, it cannot be directly optimized with gradient-descent-based optimization techniques.
To overcome this limitation, we replace the edge set $E$ with weighted adjacency $\mathbf{P}$, where $p_{ij}$ is interpreted as the probability of $(v_{i},v_{j})\in E$, which also reflects how beneficial their interaction is. Notably, we learn different graph structures $\mathbf{P}^{(k)}$ at each $k$-th layer. Compared with using a fixed graph at each layer, we have more efficiency and flexibility in enumerating the beneficial higher-order feature interaction by this means. More specifically, by using a fixed graph structure at each layer, we can only obtain a fixed set of feature interactions. However, with the adaptive learned graph structure at each layer, the model is capable of modelling any potential feature interactions.
@@ 4.2.1 Metric Function
We aim to design a metric function between each pair of feature interactions to measure whether they are beneficial. Formally, the weight $p_{ij}$ of an edge $(v_{i},v_{j})$ is computed via a metric function $f_{s}(\mathbf{e}_{i},\mathbf{e}_{j})$. Here, we adopt a neural matrix factorization (NMF) He et al. (2017) based function to estimate the edge weight. Formally, a multi-layer perception (MLP) with one hidden layer is used to transform the element-wise product of these feature vectors to a scalar:
@@
where $\mathbf{W}^{s}_{1}$, $\mathbf{W}^{s}_{2}$, $\mathbf{b}^{s}_{1}$, and $\mathbf{b}^{s}_{2}$ are the parameters of the MLP. $\delta(\cdot)$ and $\sigma(\cdot)$ are ReLU and sigmoid activation functions, respectively. It should be noted that $f_{s}$ is invariant to the order of its input, i.e., $f_{s}(\mathbf{e}_{i},\mathbf{e}_{j})=f_{s}(\mathbf{e}_{j},\mathbf{e}_{i})$. Therefore, the estimated edge weights are identical for the same pair of nodes. Such continuous modelling of the graph structure enables backpropagation of the gradients. Since we do not have a ground-truth graph structure, the gradients come from the errors between the model output with and the target.
Intuitively, we treat the element-wise product of each pair of feature embeddings as a term and estimate its weight using the MLP. One can also choose the Euclidean distance Kazi et al. (2020) or other distance metrics.
@@ 4.2.2 Graph Sampling
From the estimated edge weighted matrix $\mathbf{P}^{(k)}$ at each layer, we then sample the beneficial feature interactions to sample the neighborhood for each feature field. In this work, we uniformly sample a fixed-size set of neighbors. For each feature node $v_{i}$ at the $k$-th layer, we select $m_{k}$ edges according to the first $m_{k}$ elements of $\mathbf{P}^{(k)}[i,:]$, which can be illustrated as follows:
@@
where $\mathbf{P}^{(k)}[i,:]$ denotes the $i$-th column of matrix $\mathbf{P}^{(k)}$ at the $k$-th layer, and $\mathbf{P}^{(k)}[i,-\text{idx}_{i}]$ contains a subset of columns of $\mathbf{P}^{(k)}$ that are not indexed by $\text{idx}_{i}$. $\text{argtop}_{m_{k}}$ is an operator that selects the $m_{k}$-most important nodes for query node $i$. We only retain these $m_{k}$ feature nodes, and the others are masked. Thus the neighborhood set of node $v_{j}$ is defined as:
@@
Practically speaking, we found that our approach could achieve high performance when $k=3$, and $m_{1}$ equals the number of feature fields, which means that in the first layer, we model all pairs of feature interactions.
It is worth mentioning that we also tried to set a threshold to select the edges in the graph, i.e., setting a minimum value for the edge probability of cutting edges off. However, the performance is not as good as that of using a fixed-degree graph. This is reasonable because the edge weights of different nodes’ neighbors are at different scales. Setting a single threshold on all the nodes will lead to the situation in which the numbers of nodes’ neighbors vary greatly. Some nodes will have barely any adjacent nodes after cutting off, while some may still have many.
@@ 4.3 Interaction Aggregation
Since we have selected the beneficial feature interactions, or in other words, learned the graph structure, we perform the interaction (neighborhood) aggregation operation to update the feature representations.
@@
This indicates the importance of the interactions between feature $v_{i}$ and feature $v_{j}$.
Note that we only compute $c_{ij}$ for nodes $j\in\mathcal{N}_{i}$, where $\mathcal{N}_{i}$ denotes the neighborhood of node $v_{i}$, which is also the set of features whose interactions with $v_{i}$ are beneficial. To make coefficients easily comparable across different feature nodes, we normalized them across all choices of $j$ using a softmax function:
@@
Once the normalized attention coefficients are obtained, we compute the linear combination of these feature interactions with nonlinearity as the updated feature representations:
@@
where $\alpha_{ij}$ measures the attention coefficients of each feature interaction between feature $i$ and $j$, while $p_{ij}$ represents the probability of this feature interaction being beneficial. The attention coefficient $\alpha_{ij}$ is calculated by the soft attention mechanism, while $p_{ij}$ is calculated by the hard attention mechanism. By multiplying them together, we control the information of selected feature interactions and make the parameters in the interaction selection component trainable with gradient back-propagation.
To capture the diverse polysemy of feature interactions in different semantic subspaces Li et al. (2020) and stabilize the learning process Vaswani et al. (2017); Veličković et al. (2018), we extend our mechanism to employ multi-head attention. Specifically, $H$ independent attention mechanisms update Equation 12, and then these features are concatenated, resulting in the following output feature representation:
@@
where $\parallel$ denotes the concatenation, $\alpha_{ij}^{h}$ is the normalized attention coefficient computed by the $h$-th attention mechanism, and $\mathbf{W}^{h}_{a}$ is the corresponding linear transformation matrix. One can also choose to employ average pooling to update the feature representations:
@@ 4.4 Prediction and Optimization
The output of each $k$-th layer, is a set of $n$ feature representation vectors encoding feature interactions of order up to $k$, namely $\left\{\mathbf{e}^{(k)}_{1},\mathbf{e}^{(k)}_{2},\dots,\mathbf{e}^{(k)}_{n}\right\}$. Since the representations obtained in different layers encode the interactions of different orders, they have different contributions to the final prediction. As such, we concatenate them to constitute the final representation of each feature Wang et al. (2019):
@@
Finally, we employ average pooling on the vectors of all features to obtain a graph-level output and use a projection vector $\mathbf{p}$ to make the final prediction:
@@
where $N$ is the total number of training instances, and $\sigma$ denotes the sigmoid function. $y_{i}$ and $\hat{y}_{i}$ denote the label of instance $i$ and the prediction of GraphFM, respectively. The model parameters are updated using AdamKingma and Ba (2014).
@@ 5 Experiments
Table 1: Statistics of the evaluation datasets.
@@ 5.1 Experimental Settings
Our experiments are conducted on three real-world datasets, two CTR benchmark datasets, and one recommender system dataset. Details of these datasets are listed in Table 1. The data preparation follows the strategy in Tian et al. (2023). We randomly split all the instances at 8:1:1 for training, validation, and testing. We adopt the two most popular metrics, AUC and Logloss to measure the probability that one prediction diverges from the ground truth.
@@ 5.1.1 Baselines
We compare GraphFM with four classes of state-of-the-art methods:(A) the linear approach that only uses individual features; (B) FM-based methods that consider second-order feature interactions; (C) DNN-based methods that model high-order feature interactions; and (D) aggregation-based methods that update features’ representation and model their high-order feature interactions via an aggregation strategy.
@@
  • LR (A) refers to linear/logistics regression, which can only model linear interactions.
  • FM Rendle (2012)  (B) is the official FM implementation, which can only model second-order interactions.
  • AFM Xiao et al. (2017) (B) is an extension of FM that considers the weights of different second-order feature interactions by using attention mechanisms.
  • AFN Cheng et al. (2020)  (B) learns arbitrary-order feature interactions adaptively from data, instead of explicitly modelling all the cross features within a fixed maximum order.
@@
  • NFM He and Chua (2017) (C) devises a bi-interaction layer to model second-order interactions and uses a DNN to introduce nonlinearity and model high-order interactions.
@@
  • DeepCrossing Shan et al. (2016)  (C) utilizes a DNN with residual connections to model non-linear feature interactions in an implicit manner.
  • CrossNet Wang et al. (2017) (C) is the core of the Deep&Cross model, which models feature interactions explicitly by taking the outer product of the concatenated feature vector at the bit-wise level.
  • xDeepFM Lian et al. (2018)  (C) takes the outer product of the stacked feature matrix at a vector-wise level to explicitly model feature interactions. ANNs can also be combined with DNNs, which model implicit and explicit interactions simultaneously.
  • DCNV2 Wang et al. (2021)  (C) utilizes a cross network from the DCN to learn explicit and bounded-degree cross features.
  • AutoInt Song et al. (2019) (D) uses a self-attention network to learn high-order feature interactions explicitly. It can also be seen as performing the multi-head attention mechanism Vaswani et al. (2017) on a fully-connected graph.
@@
  • InterHAt Li et al. (2020) (D) utilizes an attention mechanism to aggregate features, which are then multiplied by the original features to produce higher-order feature interactions.